OpenStudy (anonymous):

What's the remainder when you divide 16^103 by 11?

OpenStudy (anonymous):

just do 16^103 on your scientific calculator and then divide the number by 11

OpenStudy (anonymous):

Error...

OpenStudy (anonymous):

1.458824

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

The remainder must be an integer.

OpenStudy (anonymous):

less than 11

OpenStudy (anonymous):

I'm really rusty with my number theory, but I think this is how you do it: 16^103=(16^)(16^5)(16^2)(16^1)(16^0). I'm going to use = for "congruent" 16^0=1mod11 16^1=5mod11 16^2=5^2mod11=3mod11 16^5=3^3mod11=5mod11 16^6=5^2mod11=3mod11 So you get 16^103=1*5*3*5*3=5mod11 I may have made a mistake in there, but I think the steps are correct. It's called successive squaring.

OpenStudy (anonymous):

Right idea, wrong answer:--it's 4

OpenStudy (anonymous):

Shoot, where did I mess up?

OpenStudy (anonymous):

Not sure, I went a different way, do you want to see it?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

Oh, I see what I did wrong. I messed up on the powers of 5 and 3

OpenStudy (anonymous):

16 = 5^103 = (5^10)^10 * 5^3 = 125 = 4 (all mod11)

OpenStudy (anonymous):

Ah, we used the same method, but you were smarter about it.

OpenStudy (anonymous):

Yes, u have flt a^p-1 = 1 (mod p)

OpenStudy (anonymous):

I always forget about Fermat...

OpenStudy (anonymous):

He'd be terribly upset to hear that...:-)

OpenStudy (anonymous):

At least I'm not confined to writing in the margin

OpenStudy (anonymous):

I had Goldbach on a serviette, the waiter took it away....

OpenStudy (anonymous):

Haha, nice

OpenStudy (anonymous):

Lemme try this again: 16^103=(16^64)(16^32)(16^4)(16^2)(16^1) *16^1=5mod11 *16^2=3mod11 *16^4=9mod11 16^8=4mod11 16^16=5mod11 *16^32=3mod11 *16^64=9mod11 16^103=5*3*9*3*9=3645=4mod11 Much better. Maybe I should have consulted a book before I wrote anything...

OpenStudy (anonymous):

Since we know 16^10 is congruent to 1 mod 11, we should use the division algorithm on 103 to get: \[16^{103} = 16^{10(10)+3} = (16^{10})^{10}(16^{3}) \equiv (1^{10})(16^{3}) \equiv 16^{3}\] but 16 is congruent to 5 mod 11, so we have: \[16^{3} \equiv 5^{3} \equiv 125 \equiv 4 \mod 11\]